OpenStudy (anonymous):

2x^3-x^2y-11xy^2+10y^3 / x-2y

OpenStudy (anonymous):

multiply the first equation by -2 to get: -2x - 2y = -6 if you add the first and second equations together: -2x - 2y = -6 + 2x + 2y = 10 ________________ 0x + 0x = 4 since that's impossible, the answer is c: No solutions. Another way to solve it is to substitute: y = 3 - x so 2x + 2(3 - x) = 10 2x + 6 - 2x = 10 6 = 10

OpenStudy (anonymous):

this is how @KE8717504

OpenStudy (anonymous):

so what would i enter in the little box that says quotient?

OpenStudy (anonymous):

?